OpenStudy (anonymous):

Solid Geometry, A vertical pole BP stands at one corner of a horizontal rectangular field. If AB=10m, AD=5m and the angle of elevation of P from A is 22 degrees. Calculate : (a) the height of the pole. (b) the angle of elevation of P from C. (c) the length of a diagonal of the rectangle ABCD. (d) the angle of elevation of P from D.

OpenStudy (anonymous):



OpenStudy (anonymous):

a) tan22 = BP / 10m tan22 (10m) =BP b) BP/5m = tanC c) \[\sqrt{(10^{2}) +(5^{2})} = diagonal of the rectangle\] = sci-cal can help but i forgot mine .. not quite sure with my solutions though

OpenStudy (anonymous):

indeed it's correct. I checked that with my teacher just today & what you wrote is correct. Thanks a million for your helpful comment. :)

OpenStudy (anonymous):